Please answer all questions correctly! And clear thank u
Question 26 0.5 pts A compound containing carbon and hydrogen underwent combustion to produce 4.50 g carbon dioxide and 1.84 g of water. What is the empirical formula of the compound? CH
Please answer all questions correctly! And clear t
Answer #1

The general combustion reaction is

C_xH_y + (x + y/4)O_2 \rightarrow xCO_2 + \frac{y}{2}H_2O

Molar mass of CO2 = 12 + 2 * 16 = 44 gm/mol

Number of moles of CO2 = 5.40/44 = 4.50/44 = 0.1022 moles

Number of moles of H2O = 1.84/18 = 0.1022 moles

Hence we get x=y/2 or 2x=y

Therefore the empirical formula will be CH2

Hence the correct answer is Option C

Q27)

Mass of Carbon in Compound = 54.4/100 * 88.11 = 47.93 gms

Mass of Hydrogen in Compound - 9.2/100 * 88.11 = 8.10612 grams

Mass of Oxygen in the compound = 36.3/100 * 88.11 = 31.98 grams

Hence the formula will be C4H8O2

Hence the correct answer is Option D